# Motion Artifact Removal in Pixel-Frequency Domain via Alternate Masks and Diffusion Model

### v1

@@ Introduction
We select three classes of approaches for comparison. The first is supervised approach, Pix2pix (Isola et al. 2017), which uses paired images for training. The second is unsupervised approaches, CycleGAN (Zhu et al. 2017) and UDDN (Wu et al. 2023), we only use simulated motion-corrupted images for training, as training with real images often leads to unstable training (Oh et al. 2023). The third is unsupervised approaches based on diffusion model, DR2 (Wang et al. 2023) and GDP (Fei et al. 2023), both are novel approaches for image recovery. The training parameters are set according to the parameters specified in the original papers, respectively. More information on training and evaluation settings can be found in supplementary material A.3.

### v2

@@ Introduction
We select three classes of approaches for comparison. The first is supervised approach, Pix2pix (Isola et al. 2017), which uses paired images for training. The second is unsupervised approaches, CycleGAN (Zhu et al. 2017) and UDDN (Wu et al. 2023), we only use simulated motion-corrupted images for training, as training with real images often leads to unstable training (Oh et al. 2023). The third is unsupervised approaches based on diffusion model, DR2 (Wang et al. 2023) and GDP (Fei et al. 2023), both are novel approaches for image recovery. The training parameters are set according to the original paper, respectively. For more information on training and evaluation settings, including radiologist evaluation, see supplementary material A.3.
@@ B.3 Comparison of Visualization Results on Ablation Study of Pixel-Frequency Domain
B.4 Experiment on Cutoff
We perform the following experiment on the cutoff frequency of the filter.
Cutoff PSNR SSIM Lpips FSIM VIF VSI GMSD
$\pi/5$ 25.89 0.711 0.283 0.826 0.306 0.950 0.112
$\pi/10$ 27.60 0.829 0.147 0.896 0.426 0.970 0.061
$\pi/20$ 26.34 0.779 0.180 0.867 0.343 0.962 0.085
Table 7: Experiment on the cutoff frequency of the filter.
If the cutoff is too small, the useful guided information is too small to recover the correct texture, if the cutoff is too large, the guided information contains too many artifact information, both of them lead to bad results. Therefore, it is necessary to correctly select the cutoff frequency in clinical diagnosis.
B.5 Experiment on Grid Size
We have used a checkerboard mask with the grid size of 16. We perform an experiment on the grid size. The results are as follows.
Size PSNR SSIM Lpips FSIM VIF VSI GMSD
1 27.57 0.829 0.153 0.896 0.423 0.970 0.062
4 27.57 0.829 0.153 0.896 0.423 0.970 0.063
16 27.60 0.829 0.147 0.896 0.426 0.970 0.061
64 27.42 0.827 0.154 0.895 0.420 0.969 0.064
128 27.24 0.824 0.155 0.893 0.416 0.969 0.065
Table 8: Experiment on the grid size of checkerboard mask.
Large grid size destroys the distribution of useful information and small grid size does not significantly destroy the distribution of artifact, both of them lead to poorer results, the former is worse.